OpenStudy (anonymous):

help will give medal The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket: 132 = one half x + 2(x − 9) What is the regular price of a child's ticket? (5 points) $60 $57 $51 $42

OpenStudy (anonymous):

That equation is correct. 132 = 1/2x + 2(x - 9) Solve for x to get the regular price of an adult ticket. 132 = 1/2x + 2x - 18 150 = 2.5 x 60 = x Subtract 9 to get the regular price of a child's ticket.

OpenStudy (anonymous):

oh I see thank you so much

OpenStudy (anonymous):

Your welcome love :)